Summary

Software debugging and system check after software merged.

Team: Fabio and Alfio.

Activity:

* Use git to get the same "merged" version. No problems presented. OptLoopDiagnostic code reinstalled.
* Adsdec power on showed several issues:
  • Actuator 569 wron position.
  • IDL Dust detected in the mirror.After power cycle wasn't detected again. To be investigated.
  • Loop closes but no slopes are received. After power cycle Bcu2K was OK.
  • Shell RIP during closed loop - the fast diagnostic log "Timeout - will emergency react now". This usually happens when the adsec BCUs stop sending telemetry. Another adsec power cycle.
* Close loop showed issues to get working the optical gain (not demodulating), this means that the information from the secondary positions and commands are zeros when a TN is saved. The process was restarted several times and it showed the same behaviour.

The arbitrator should shows a everytimt ethe loop is clopse a line information like:
MasterDiagnostic_33276|INF|      8803|2021-01-22 17:43:44.642791|            MAIN > Setting optical loop CLOSED

After a new presaet was sent from the AOSGUI, evertyhting worked again. ??

Arcetri is investigating this behaviour.

Instrument release to LBTO.
power OFF.
ARGOS CAL UNIT retracted.

12:10 PM Handover for telescope activities.

-- JenniferPower - 25 Jan 2021
Topic revision: r1 - 25 Jan 2021, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback